OpenStudy (anonymous):

2. The height of one solid limestone square pyramid is 24 m. A similar solid limestone square pyramid has a height of 30 m. The volume of the larger pyramid is 16,000 m3. Determine each of the following, showing all your work and reasoning.

OpenStudy (anonymous):

a) The scale factor of the smaller pyramid to the larger pyramid in simplest form. 24/30 = 4/5/4:5(going based on height) b) The ratio of the area of the base of the smaller pyramid to the larger pyramid. (4/5)^2 = c) Ratio of the volume of the smaller pyramid to the larger. d) The volume of the smaller pyramid.